# PlasmoData.jl – A Julia Framework for Modeling and Analyzing Complex Data as Graphs

### v1

@@ 19                    EC = get_EC(filtered_graph)
27        # Create array for storing solutions
28        ECs = zeros(size(data, 1) - 6)
29        num_max_cliques = zeros(size(data, 1) - 6)
30        num_communities = zeros(size(data, 1) - 6)
@@ 48            ECs[i], num_max_cliques[i], num_communities[i] = find_smallest_filtered_graph(sym_graph)
With the filtered DataGraph defined, we can consider several topological metrics for analyzing our data. The goal of de Souza and co-workers was to identify ways of “fingerprinting” when disease outbreaks were occurring. They chose the EC as their primary indicator, and they showed how the EC was able to help indicate when outbreaks were occurring, especially for larger $d_{max}$ values. This may be because the higher $d_{max}$ values indicate when there are several nodes (districts) are closely correlated. Consequently, we consider the EC of the graphs being formed, and then we also consider the size of cliques in the graph and the number of communities in the graph. In graph theory, a $k$-clique is a set of $k$ nodes which are connected to all other nodes in the clique. We hypothesize that a possible fingerprint for outbreaks may be when several nodes (districts) are closely correlated. Thus, we compute the EC for comparison with de Souza and co-workers, we compute the total number of maximal cliques in the graph (i.e, the total number of $k_{max}$-cliques in the graph, where $k_{max}$ is the largest integer value for which there is a non-empty set of nodes forming a $k_{max}$-clique in the graph) and we compute the number of communities in the graph (found by clique percolation[46], with $k$ = 25; see Lines 18 - 23). The results of these computations, as well as the average daily new cases during each 7 day moving window, are shown in Figure 14.
@@
Figure 14: Results for TDA of dengue data for 93 districts in Recife, Brazil[13]. For each 7-day moving time window, an edge-weighted DataGraph was formed. Results presented include the average number of new cases per day for the moving time window (a), the EC for the DataGraph (b), the number of maximal cliques in the DataGraph (c), and the number of communities in the DataGraph using clique percolation ($k=25$)[46].
@@
The results shown in Figure 14 are consistent with the results of de Souza and co-workers. As mentioned above, the EC we compute in the above code is for a graph, while de Souza and co-workers computed the EC for CW-complexes, where their computed EC was a function of $d_{max}$. They used various $d_{max}$ values up to $7$, and found that the noise of the EC decreased with increasing $d_{max}$. In this case, the EC in Figure 14 does contain significant noise (visually similar to that seen for $d_{max}=2$ and $d_{max}=3$ of de Souza and co-workers’ work). However, the trends are similar, and we note that it is computationally much easier to compute the EC for a graph (i.e., number of nodes minus number of edges) than to compute the EC for a CW-complex with larger $d_{max}$ (where each $d_{max}$-cell may have to be determined). In addition, the number of maximal cliques (Figure 14c) and the number of communities (Figure 14d) also correlate to the average number of new cases per day, and they appear slightly less noisy than the EC. A detailed analysis of these additional indicators as fingerprints for disease outbreaks is outside the scope of this work; instead, our purpose in showing these results is to highlight how the DataGraph abstraction and how PlasmoData.jl facilitates different TDA approaches.
@@ 4.3 Connectivity Analysis
Different types of nodes have specific data attached to them. For example, nodes can contain supply limitations, demands, costs, and carbon emissions, and these will be saved as node data. This problem also contains connections between raw materials and technologies, technologies and intermediates, intermediates and technologies, and technologies and products which can all be represented by directed edges, forming a directed graph, with 57 total edges. In addition, we also have flows that are represented as edge data. These flows are dependent on the connections in the graph and on the data stored in the nodes (e.g., it depends on the cost of technologies or the maximum available supply of the products).
The DataGraph contains the set of 54 nodes and 57 directed edges discussed above, which we notate here as $\mathcal{N}$ and $\mathcal{E}$. There are four subsets of nodes, $\mathcal{N}_{raw},\mathcal{N}_{prod},\mathcal{N}_{int}$, and $\mathcal{N}_{tech}$ for the sets of raw materials, products, intermediates, and technologies, respectively, such that $\mathcal{N}=\mathcal{N}_{raw}\cup\mathcal{N}_{prod}\cup\mathcal{N}_{int}\cup\mathcal{N}_{tech}$. Each node stores data under different attributes indicating whether the node is a raw material ($a_{raw}$), product ($a_{prod}$), intermediate ($a_{int}$), or technology ($a_{tech}$) and data under different attributes for the raw material cost ($a_{cost}$), the CO2 emissions per weight of material ($a_{CO_{2}}$), maximum raw material supply ($a_{sup}$), and maximum production demand ($a_{dem}$). Each edge has data for the optimal flow of material along the edge under attribute $a_{flow}$.
@@
where $A$ is an asymmetric matrix ($\mathcal{DG}$ is directed) matching the connectivity defined by $\mathcal{E}$.
Here, $\alpha_{cost}(n)$ is the raw material cost for node $n$, $\alpha_{CO_{2}}(n)$ is the CO2 emissions per weight of material for node $n$, $\alpha_{sup}(n)$ is the maximum raw material supply for node $n$, $\alpha_{dem}(n)$ is the maximum production demand for node $n$, and $\alpha_{flow}(e)$ is the optimal flow on edge $e$ (this value is computed separately by solving an optimization problem that uses the above data).
@@
1        using PlasmoData, Graphs, DelimitedFiles, DataGraphPlots
@@
3        # Read in data
4        raw_data  = readdlm("rawmaterial_data_array.csv", ’,’)
5        prod_data = readdlm("product_data_array.csv", ’,’)
6        int_data  = readdlm("intermediates_data_array.csv", ’,’)
7        tech_data = readdlm("technology_data_array.csv", ’,’)
8        edge_data = readdlm("edge_data.csv", ’,’)
@@
10        # Define DataDiGraph
11        dg = DataDiGraph()
@@
13        # Add nodes and node data
14        for i in 1:size(raw_data, 1)
@@
22        for i in 1:size(prod_data, 1)
@@
33        for i in 1:size(tech_data, 1)
@@
40        # Add edges and edge data
41        for i in 1:size(edge_data, 1)
@@
43            DataGraphs.add_edge!(dg, edge)
44            DataGraphs.add_edge_data!(dg, edge, edge_data[i, 3], "Optimal Flow")
45        end
@@
47        DataGraphPlots.plot_graph(dg, dag_positions = true, nlabels = dg.nodes)
@@ 4.3.1 Visualization
One of the potential benefits of these graph representations is the ability to visualize the structure of the problem. In this example, $\mathcal{DG}$ is a directed acyclic graph (DAG), which is a common structure in pathway and supply chain problems. Because it is a DAG, the graph can be presented in different layers, and different algorithms have been proposed for identifying these layers and the positions of nodes within the visualization [81, 82]. The above technology problem is shown in Figure 15 using the algorithm of Zarate and co-workers [82] to determine node positions (as implemented within the Julia package LayeredLayouts.jl [83]). This visualization is automated through DataGraphPlots.jl. In this case, Figure 15 (without node coloring) can be easily constructed with just Line 47 of Code Snippet 10 (abbreviation meanings are available in the supporting information).
Figure 15: Directed Graph of a layered technology pathway example
The ability to visualize these structures can be a significant resource in evaluating the data. For example, it can make clear what some of the paths are or what some of the most essential nodes are. For instance, from Figure 15, it is clear that the node ”Ethylene” has more connections than any other node and is directly connected to several products. In addition, DataGraphPlots.jl provides the ability to highlight paths within a graph. For example, by calling the function plot_graph_path with the source (”Corn Stover”) and destination (”Nylon”) nodes of the path, Figure 16 can be generated. Visualizations for DAGs or for layered graphs are often not implemented directly in graph plotting tools, and other tools for visualization (e.g., yFiles [84]) can require subscriptions, so this ability for visualization is a powerful capability of PlasmoData.jl.
Figure 16: Directed graph with a pathway between ”Corn Stover” and ”Nylon” highlighted.
@@
Code Snippet 11: Code for constructing a Sankey diagram from a DataDiGraph
1        using SankeyPlots
@@
3        # Define src and dst for edges
4        src = [i for (i, j) in dg.edges]
5        dst = [j for (i, j) in dg.edges]
6        weights = get_edge_data(dg, "Optimal Flow")
@@
8        # Plot sankey diagram
9        sankey(src, dst, weights;
@@
14        )
Figure 17: Sankey diagram of a technology pathway example.
@@ 4.3.2 Connectivity Measures
1        # Compute metrics for raw materials
2        for r in raw_data[:, 1]
@@
14        # Compute metrics for products
15        for p in prod_data[:, 1]
@@
Table 1: Metrics on the technology graph with the number of connected products (Prod) to each raw material and the number of downstream nodes (Down) for each raw material. The nodes Ethylene, Ter Acid, and C.-hexane were each individually removed from the original graph and metrics recomputed.
@@
Table 2: Metrics on the technology graph with the number of connected raw materials (Raw) to each product and the number of upstream nodes (Up) for each product. The nodes Ethylene, Ter Acid, and C.-hexane were each individually removed from the original graph and metrics recomputed.
@@
Figure 18: Directed graph of a layered technology pathway example with nodes sized by the number of downstream nodes
@@ 5 Conclusions and Future Work
There are a couple of areas that we would like to focus on in future work. First, there are more applications we would like to explore. Using graphs to represent molecules has garnered significant interest and has been applied in many studies and we would like to streamline molecular representations within PlasmoData.jl, such as enabling construction of the graph from a SMILES (Simplified Molecular Input Line Entry System) string, and explore how PlasmoData.jl can enhance this analysis. We are also interested in other applications of the pathway analysis performed above. Additionally, we would like to expand the capabilities of PlasmoData.jl. There are other TDA metrics and path functions that could be implemented, and there are topological descriptors that could be approximated with graphs. For example, in the work by Jiang and co-workers[69], they use a marching-square algorithm to approximate the Minkowski functionals of a filtered image. A marching-square algorithm could be implemented for an image represented by a graph by iterating through the nodes and analyzing the connections to the adjacent nodes. This could lead to more in-depth image analysis. Also, we would like to expand the integration of PlasmoData.jl with GNN packages.
In addition, while graphs can be powerful tools for modeling data, they do have some limitations. Nodes and edges do not have any notion of placement in space. Oftentimes, the systems being represented by graphs have a fixed location (e.g., supply chains may have a fixed production location) which can significantly influence the system. Furthermore, graphs only capture connectivity between nodes; in many instances, it may be more accurate to include a link between more than two nodes. In these cases, hypergraphs (or other representations) may be more appropriate. Finally, graphs are inherently 2D objects; in future work, we would like to explore generalizations of hypergraphs and simplicial complexes and how to capture more complex attributes for these objects, such as spatial locations.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research (ASCR) under contract DE-0002722 and in part by the National Science Foundation under award CBET-2315963.

### v2

@@ 19 EC = get_EC(filtered_graph)
27# Create array for storing solutions
28ECs = zeros(size(data, 1) - 6)
29num_max_cliques = zeros(size(data, 1) - 6)
30num_communities = zeros(size(data, 1) - 6)
@@ 48 ECs[i], num_max_cliques[i], num_communities[i] = find_smallest_filtered_graph(sym_graph)
With the filtered DataGraph defined, we can consider several topological metrics for analyzing our data. The goal of de Souza and co-workers [13] was to identify ways of “fingerprinting” when disease outbreaks were occurring. They chose the EC as their primary indicator, and they showed how the EC was able to help indicate when outbreaks were occurring, especially for larger $d_{max}$ values. This may be because the higher $d_{max}$ values indicate when there are several nodes (districts) are closely correlated. Consequently, we consider the EC of the graphs being formed, and then we also consider the size of cliques in the graph and the number of communities in the graph. In graph theory, a $k$-clique is a set of $k$ nodes which are connected to all other nodes in the clique. We hypothesize that a possible fingerprint for outbreaks may be when several nodes (districts) are closely correlated. Thus, we compute the EC for comparison with de Souza and co-workers [13], we compute the total number of maximal cliques in the graph (i.e, the total number of $k_{max}$-cliques in the graph, where $k_{max}$ is the largest integer value for which there is a nonempty set of nodes forming a $k_{max}$-clique in the graph) and we compute the number of communities in the graph (found by clique percolation [46], with $k$ = 25; see Lines 18 - 23). The results of these computations, as well as the average daily new cases during each 7 day moving window, are shown in Figure 14.
@@
Figure 14: Results for TDA of dengue data for 93 districts in Recife, Brazil [13]. For each 7-day moving time window, an edge-weighted DataGraph was formed. Results presented include the average number of new cases per day for the moving time window (a), the EC for the DataGraph (b), the number of maximal cliques in the DataGraph (c), and the number of communities in the DataGraph using clique percolation ($k=25$) [46]
@@
The results shown in Figure 14 are consistent with the results of de Souza and co-workers [13]. As mentioned above, the EC we compute in the above code is for a graph, while de Souza and co-workers [13] computed the EC for CW-complexes, where their computed EC was a function of $d_{max}$. They used various $d_{max}$ values up to $7$, and found that the noise of the EC decreased with increasing $d_{max}$. In this case, the EC in Figure 14 does contain significant noise (visually similar to that seen for $d_{max}=2$ and $d_{max}=3$ of de Souza and co-workers’ [13] work). However, the trends are similar, and we note that it is computationally much easier to compute the EC for a graph (i.e., number of nodes minus number of edges) than to compute the EC for a CW-complex with larger $d_{max}$ (where each $d_{max}$-cell may have to be determined). In addition, the number of maximal cliques (Figure 14c) and the number of communities (Figure 14d) also correlate to the average number of new cases per day, and they appear slightly less noisy than the EC. A detailed analysis of these additional indicators as fingerprints for disease outbreaks is outside the scope of this work; instead, our purpose in showing these results is to highlight how the DataGraph abstraction and how PlasmoData.jl facilitates different TDA approaches.
@@ 4.3 Connectivity Analysis
Within the above technology pathway, different types of nodes have specific data attached to them. For example, nodes can contain supply limitations, demands, costs, and carbon emissions, and these will be saved as node data. This problem also contains connections between raw materials and technologies, technologies and intermediates, intermediates and technologies, and technologies and products which can all be represented by directed edges, forming a directed graph, with 57 total edges. In addition, we also have flows that are represented as edge data. These flows are dependent on the connections in the graph and on the data stored in the nodes (e.g., it depends on the cost of technologies or the maximum available supply of the products).
The DataGraph for this technology pathway is a heterogenous graph and contains multiple attributes for node and edge data. The DataGraph contains the set of 54 nodes and 57 directed edges discussed above, which we notate here as $\mathcal{N}$ and $\mathcal{E}$. There are four subsets of nodes, $\mathcal{N}_{raw},\mathcal{N}_{prod},\mathcal{N}_{int}$, and $\mathcal{N}_{tech}$ for the sets of raw materials, products, intermediates, and technologies, respectively, such that $\mathcal{N}=\mathcal{N}_{raw}\cup\mathcal{N}_{prod}\cup\mathcal{N}_{int}\cup\mathcal{N}_{tech}$. Each node stores data under different attributes indicating whether the node is a raw material ($a_{raw}$), product ($a_{prod}$), intermediate ($a_{int}$), or technology ($a_{tech}$) and data under different attributes for the raw material cost ($a_{cost}$), the CO2 emissions per weight of material ($a_{CO_{2}}$), maximum raw material supply ($a_{sup}$), and maximum production demand ($a_{dem}$). Each edge has data for the optimal flow of material along the edge under attribute $a_{flow}$.
@@
where $A$ is an asymmetric matrix ($\mathcal{DG}$ is directed) matching the connectivity defined by $\mathcal{E}$. Here, $\alpha_{cost}(n)$ is the raw material cost for node $n$, $\alpha_{CO_{2}}(n)$ is the CO2 emissions per weight of material for node $n$, $\alpha_{sup}(n)$ is the maximum raw material supply for node $n$, $\alpha_{dem}(n)$ is the maximum production demand for node $n$, and $\alpha_{flow}(e)$ is the optimal flow on edge $e$ (this value is computed separately by solving an optimization problem that uses the above data).
@@
1using PlasmoData, Graphs, DelimitedFiles, PlasmoDataPlots
@@
3# Read in data
4raw_data = readdlm("rawmaterial_data_array.csv", ’,’)
5prod_data = readdlm("product_data_array.csv", ’,’)
6int_data = readdlm("intermediates_data_array.csv", ’,’)
7tech_data = readdlm("technology_data_array.csv", ’,’)
8edge_data = readdlm("edge_data.csv", ’,’)
@@
10# Define DataDiGraph
11dg = DataDiGraph()
@@
13# Add nodes and node data
14for i in 1:size(raw_data, 1)
@@
22for i in 1:size(prod_data, 1)
@@
33for i in 1:size(tech_data, 1)
@@
40# Add edges and edge data
41for i in 1:size(edge_data, 1)
@@
43 PlasmoData.add_edge!(dg, edge)
44 PlasmoData.add_edge_data!(dg, edge, edge_data[i, 3], "Optimal Flow")
45end
@@
47plot_graph(dg, dag_positions = true, nlabels = dg.nodes)
@@ 4.3.1 Visualization
One of the potential benefits of these graph representations is the ability to visualize the structure of the problem. In this example, $\mathcal{DG}$ is a directed acyclic graph (DAG), which is a common structure in pathway and supply chain problems. Because it is a DAG, the graph can be presented in different layers, and different algorithms have been proposed for identifying these layers and the positions of nodes within the visualization [81, 82]. The above technology problem is shown in Figure 15 using the algorithm of Zarate and co-workers [82] to determine node positions (as implemented within the Julia package LayeredLayouts.jl [83]). This visualization is automated through PlasmoDataPlots.jl. In this case, Figure 15 (without node coloring) can be easily constructed with just Line 47 of Code Snippet 10 (abbreviation meanings are available in the supporting information).
Figure 15: Directed graph of a layered technology pathway example; created using PlasmoDataPlots.jl
The ability to visualize these structures can be a significant resource in evaluating the data. For example, it can make clear what some of the paths are or what some of the most essential nodes are. For instance, from Figure 15, it is clear that the node ”Ethylene” has more connections than any other node and is directly connected to several products. In addition, PlasmoDataPlots.jl provides the ability to highlight paths within a graph. For example, by calling the function plot_graph_path with the source (”Corn Stover”) and destination (”Nylon”) nodes of the path, Figure 16 can be generated. Visualizations for DAGs or for layered graphs are often not implemented directly in graph plotting tools, and other tools for visualization (e.g., yFiles [84]) can require subscriptions, so this ability for visualization is a powerful capability of PlasmoData.jl.
Figure 16: Directed graph with a pathway between ”Corn Stover” and ”Nylon” highlighted, created using PlasmoDataPlots.jl
@@
Code Snippet 11: Code for constructing a Sankey diagram from a PlasmoData.DataDiGraph
1using SankeyPlots
@@
3# Define src and dst for edges
4src = [i for (i, j) in dg.edges]
5dst = [j for (i, j) in dg.edges]
6weights = get_edge_data(dg, "Optimal Flow")
@@
8# Plot sankey diagram
9sankey(src, dst, weights;
@@
14)
Figure 17: Sankey diagram of a technology pathway example
@@ 4.3.2 Connectivity Measures
1# Compute metrics for raw materials
2for r in raw_data[:, 1]
@@
14# Compute metrics for products
15for p in prod_data[:, 1]
@@
Table 1: Metrics on a technology graph with the number of connected products (Prod) to each raw material and the number of downstream nodes (Down) for each raw material. The nodes Ethylene, Ter Acid, and C.-hexane were each individually removed from the original graph and metrics recomputed
@@
Table 2: Metrics on the technology graph with the number of connected raw materials (Raw) to each product and the number of upstream nodes (Up) for each product. The nodes Ethylene, Ter Acid, and C.-hexane were each individually removed from the original graph and metrics recomputed
@@
Figure 18: Directed graph of a layered technology pathway example with nodes sized by the number of downstream nodes, created using PlasmoDataPlots.jl
@@ 5 Conclusions and Future Work
There are a couple of areas that we would like to focus on in future work. First, there are more applications we would like to explore. Using graphs to represent molecules has garnered significant interest and has been applied in many studies and we would like to streamline molecular representations within PlasmoData.jl, such as enabling construction of the graph from a SMILES (Simplified Molecular Input Line Entry System) string, and explore how PlasmoData.jl can enhance this analysis. We are also interested in other applications of the pathway analysis performed above. Additionally, we would like to expand the capabilities of PlasmoData.jl. There are other TDA metrics and path functions that could be implemented, and there are topological descriptors that could be approximated with graphs. For example, in the work by Jiang and co-workers [69], they use a marching-square algorithm to approximate the Minkowski functionals of a filtered image. A marching-square algorithm could be implemented for an image represented by a graph by iterating through the nodes and analyzing the connections to the adjacent nodes. This could lead to more in-depth image analysis. Also, we would like to expand the integration of PlasmoData.jl with GNN packages; this could include improving the interface with GNN packages and enabling different graph manipulation functions with GNN algorithms, such as using aggregation for pooling operations or message-aggregator operators.
In addition, while graphs can be powerful tools for modeling data, they do have some limitations. Nodes and edges do not have any notion of placement in space. Oftentimes, the systems being represented by graphs have a fixed location (e.g., supply chains may have a fixed production location) which can significantly influence the system. Furthermore, graphs only capture connectivity between nodes; in many instances, it may be more accurate to include a link between more than two nodes. In these cases, hypergraphs (or other representations) may be more appropriate. Finally, graphs are restricted to edges between no more than two nodes; in future work, we would like to explore generalizations of hypergraphs and simplicial complexes (which may capture more complex and higher-order relations) and how to capture more complex attributes for these objects, such as spatial locations.
@@ Acknowledgments
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research (ASCR) under contract DE-SC0023361 and in part by the National Science Foundation under award CBET-2315963.